[Brahms-l] Shift report 20050604 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat Jun 04 2005 - 23:57:56 EDT
Shifter leader was HHD
Shifters were 


1. It took a very long time for MCR to produce a lasting, stable store. Again they had to recover from a quench link interlock. An early, short store produced the runs marked with * below. They were quite short due to trouble with the scaledowns and interruptions caused by my attempts to fix the trigger 6 problem.

2. Data runs with FS@2.3A861, MRS@52.5B700:
   15200*,15201*,15204,15205,15206,15207,15208,15209,15210
   Non-TPC pedestal run: 15202
   TPC pedestal run: 15203

3. Still, the usual BB voltage problem. Furthermore trigger 6 is not taking data and FV is looking into this. 

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jun 4 23:58:12 2005

This archive was generated by hypermail 2.1.8 : Sat Jun 04 2005 - 23:58:24 EDT